IPP Progress Report for the week YYYY.MM.DD - YYYY.MM.DD

(Up to IPP Progress Reports)

Eugene Magnier

Serge Chastel

  • Czar on Monday
  • Fixed platescale value in ipp-mops ppTranslate (undefined value if no detections in the last cmf file entry)
  • Condor: good progress. Had many problem with the configuration (the default configuration is UCWS-site dependent). I can execute programs from pantasks. See https://svn.pan-starrs.ifa.hawaii.edu/repo/ipp/branches/sc_branches/pantasks_condor (see the documentation folder in it). "Development" made from the ippdor account ~/dev/pantasks_condor

Heather Flewelling

  • cleaned up 13.1 TB of old LAP databases on ipp005.
  • organize ipptopsps wikis (this is entirely my fault) and added a diagram showing how stuff is processed: [ippToPspsIndex]
  • ThreePi?.V3 minidvodb builds are finished, however, there is one failed one 1781 that I need to investigate

  • LAP.slice18 (see wiki) is being built.
    • the kepler area of sky for cam stage is built
    • the rest of slice18 - cam stage is currently building
    • skycal stage is not yet queued (was waiting on skycal to finish processing)
    • skycal stage needs some addtool tweaks before I can queue
    • have some ideas for how to monitor it which I need to implement

  • PSPS SA11 - ipptopsps loaded batches for SA11, unfortunately we have more things to fix so there will also be a SA12.
  • work with steph/haydn about SSP datastore
  • liason
    • files for daniel at jhu
    • sent louise information on PS1 images
  • aas meeting - booked travel

Mark Huber

  • MD04,05 refstacks finished, verified, distributed
  • MD01.GR0 reprocessing to V3 and nightly stacks wrap-up ~90% complete
  • lots of system kicking/monitoring, lm_sensors during faults appears to be symptom on wave1 machines and needs more investigation as to the true cause
  • re-checked WS diffim and psphot test photometry with trace setting, continue to look into cause of photometry change with trace setting

Bill Sweeney

Chris Waters

  • Czaring, mostly involving fighting hosts that were being uncooperative.
  • Added update to psphot to allow for an expanded definition of extended objects. This should help the trail fitting completeness.
  • Debugging and finishing background regeneration code. Need to finish testing of the pantasks task, but the scripting and coding are complete.
  • LAP seq_id=11 completed to 4 hours RA.